OpenStudy (anonymous):

f(x)=3x^2-12x+15 Vertex: (2,3) What is the equation of the line of symmetry: 2 What is the max/min:(2,3) Is this correct?

OpenStudy (anonymous):

For any parabola of the form \(ax^2+bx+c=0 \) the vertex will be at \( \large \left(-\frac b {2a} , -\frac {b^2-4ac}{4a}\right) \).Equation of line of symmetry will be at \(x=-\frac b {2a} \) Maximum and minimum values will be at \( \large \left(-\frac b {2a} , -\frac {b^2-4ac}{4a}\right) \) and \( \large \left(-\frac b {2a} , -\frac {b^2-4ac}{4a}\right) \) for \(a<0\) and \(a>0\) respectively.

OpenStudy (anonymous):

Ok I will write this down...this is how my teacher did it...